Solve each equation.
15. (x – 1)(x2 + 5x + 6) = 0


16. x3 - 10x2 + 16x = 0


17. (x + 2)(x2 + 3x – 40) = 0

18. x3 + 3x2 – 54x = 0

Answer by OnlyJust(1)

Solve each equation.
15. (x – 1)(x2 + 5x + 6) = 0
Simplify! Take away the brackets:
Add the values in the first brackets to the values in the second Brackets:
5x + x = 6x 6 + -1 = 5. So now you have
x2 + 6x + 5 = 0
Minus the five from the equation and that gives you
x2 + 6x = -5
u work out the rest dumbass
16. x3 - 10x2 + 16x = 0

17. (x + 2)(x2 + 3x – 40) = 0
18. x3 + 3x2 – 54x = 0